# A Match Made in Heaven? AI-driven Matching of Vulnerabilities and Security Unit Tests

### v3

@@ 4.1.1. Finding Task
Table 1 reports the performance of the six evaluated models and the three baseline approaches in their best configurations for the Finding task on $TE_{F}$ extracted from Vul4J (the other configurations are in the replication package (Iannone, Emanuele and Bui, Quang-Cuong and Scandariato, Riccardo, 2025)). The best model was UniXcoder (Guo et al., 2022), achieving an $F_{0.5}$ score of $0.73$, supported by a high precision of $0.83$. This means that it can often recognize real vulnerability-witnessing tests, with minimal false positives (just two). In other words, the tests that it flags are very likely to be security-related. The good performance of UniXcoder was also confirmed by the high MCC score of $0.61$, indicating a strong positive correlation with the ground truth (following the same interpretation of Pearson’s correlation coefficient (Pearson and Galton, 1895)). UniXcoder identified 10 out of 21 vulnerability-witnessing tests in $TE_{F}$, resulting in an unremarkable recall score of $0.48$. Still, this was the highest recall across all tested models. UniXcoder had slightly lower precision than Qwen2.5-Coder, which was $0.88$; however, this was due to a single false positive.
@@ 4.1.2. Matching Task
Table 2 reports the performance of the six evaluated models and the three baseline approaches in their best configurations for the Matching task on $TE_{M}$ extracted from Vul4J (the other configurations are in the replication package (Iannone, Emanuele and Bui, Quang-Cuong and Scandariato, Riccardo, 2025)). The best model was DeepSeek Coder (Guo et al., 2024), achieving an $F_{0.5}$ score of $0.65$, supported by a high precision of $0.75$. This means that it can often validate pairs of test cases and vulnerability descriptions, with minimal false positives (just three). In other words, the matches that it returns are very likely to be valid. This is also confirmed by the MCC score of $0.57$, indicating a strong positive correlation with the ground truth (Pearson and Galton, 1895).
Unlike the Finding task, the best model belongs to the LLM group. This DeepSeek Coder model was trained in “full-train” mode (PT-FT), which made a pre-training on $TR^{\prime}_{M}$ (the simplified Matching task) for eight epochs and then the fine-tuning on $TR_{M}$ (the regular Matching task) for five epochs. Both training datasets were augmented with SPAT (Yu et al., 2022) (ran five times). This confirmed that the model benefited from the two-phase training. Indeed, comparing this version with the configuration that only performed fine-tuning (FT) on $TR_{M}$, the $F_{0.5}$ score went from $0.58$ to $0.65$, due to an improved precision (from $0.64$ to $0.75$) and the same recall. The other LLMs, by contrast, achieved their best performance with only fine-tuning (i.e., pre-training actually lowered their scores). The effectiveness of two-phase training appears to vary with the specific model, making it an impactful factor. Furthermore, all LLMs showed consistent benefits from data augmentation. This is likely due to the inability to employ a weighted loss function as we did with the CRMs.
@@
Among the baseline approaches, SimMatch performed poorly. The SimMatchCRM flavor achieved greater precision, but at the cost of reducing the detection rate almost to zero (only one true positive and two false positives). Despite its simplicity, GrepMatch (with two required hits) achieved higher precision, i.e., $0.50$. This indicates that when we searched for two terms within the vulnerability descriptions in the test case text, we could validate 50% of the true pairs. This method appears suitable as an initial, lightweight approach for quickly matching a few test cases with vulnerabilities (i.e., the “low-hanging fruit”). Unfortunately, this method also missed many cases ($0.14$ recall).
@@ 4.2.1. Finding Task
As a result of the experimentation in RQ1, VuTeCo uses UniXcoder for the Finding task. After processing 1,105,491 test cases (from a total of 427 projects), VuTeCo flagged 319 test methods as “Security”, found in 83 projects. Among these, the manual assessment confirmed that 224 were covering security-related aspects—i.e., VuTeCo scored $0.70$ precision. The inspectors agreed on 294 ($92\%$) cases, with $0.81$ Cohen’s Kappa score, indicating a strong inter-rater agreement. Afterward, they jointly reviewed the remaining 25 cases, where they had conflicting judgments, until they reached consensus.
The precision score in the wild is not much far from the $0.83$ observed during the experiment in RQ1—indeed, a non-negligible drop was anticipated. Thus, the transfer of the Finder model in the wild mostly preserved its capability. We observed that VuTeCo was misled by test cases containing keywords commonly associated with vulnerabilities, even though these tests did not actually uncover any security issues. For instance, a test case in Apache Struts contains terms like ‘bad’, ‘pollution’, and ‘inject’—which are often found in security vocabularies—despite having a different meaning there (Apache Flink, 2025). Besides, we observed that test cases containing constant strings such as paths (e.g., URLs or file paths), shell commands, version numbers, or hashes are more likely to be incorrectly flagged. Such problems could be addressed by introducing fine-grained semantic analyses and improved keyword matching to substantially reduce the number of false positives.
@@ 4.2.2. Matching Task
As a result of the experimentation in RQ1, VuTeCo uses DeepSeek Coder for the Matching task. After processing 5,451,212 pairs of test cases and vulnerability descriptions, VuTeCo flagged 96 pairs as “Matched”, involving 35 unique tests from 19 projects. Among these, the manual assessment confirmed that 45 were correct, i.e., VuTeCo scored $0.47$ precision. The inspectors agreed on 85 ($89\%$) cases, with $0.77$ Cohen’s Kappa score, indicating a substantial inter-rater agreement. Afterward, they jointly reviewed the remaining 11 cases in which they had conflicting judgments until they reached consensus.
Unlike the Finding task, here we observed a larger drop in precision than the $0.75$ achieved during experimentation in RQ1. Many of the invalid matches were due to the similarity between the terms in the test case and the CVE description (just as in the Finding case). This decline in precision can be attributed to two main reasons: (i) The 427 projects chosen for this assessment might have no valid match at all, whereas the 51 projects used in RQ1 were guaranteed to have at least one valid match; (ii) the number of vulnerabilities for each project is higher than in RQ1, particularly for large projects like Spring Framework or Jenkins. Therefore, there is a natural distribution shift that increases the likelihood of false positives. Despite these issues, the transfer of the Matcher model in the wild was deemed acceptable, though it requires further improvements, including better prompting and safeguards against misleading terms.
In the end, the 35 tests correctly matched with the right vulnerability were added to the new dataset Test4Vul (Iannone, Emanuele and Bui, Quang-Cuong and Scandariato, Riccardo, 2025) alongside the 224 security-related tests from the Finding task, totaling 259.
Answer to RQ2. VuTeCo found 224 security-related test cases in the wild, i.e., $70\%$ of all tests returned. The false positives were primarily due to security-related terms and constant strings appearing in the test code. Then, VuTeCo returned 45 correct matches, i.e., $47\%$ of the total matches. Test code and CVE descriptions sharing a similar vocabulary led to incorrect matches.
@@ 5. Discussion
The Retrieval of Witnessing Tests. The main advantage of VuTeCo lies in its fully static nature. A dynamic assessment would require building an entire project and setting up the environment to execute all test cases, which may not always succeed due to dependency issues. The extent of the results that VuTeCo returned in the wild (RQ2) permits manual inspections to be done in a reasonable time. Indeed, VuTeCo greatly reduced the search space by four orders of magnitude, from millions to a few hundred cases. Hence, VuTeCo is designed as a lightweight assessment tool to find vulnerability-witnessing tests before proceeding to a more comprehensive dynamic assessment. Despite the positive results obtained from the experimental evaluation (RQ1), the retrieval of vulnerability-witnessing tests remains a challenging task. We further examined the errors made by VuTeCo to identify potential improvements. The main reason is attributable to the vocabulary of the test code. As observed, terms related to security and certain constant strings often lead AI models into error, likely due to their similarity to the vocabulary of the positive instances in the training sets for the Finding and Matching tasks. This problem highlighted the need for a dedicated pre-training session to enable the model to adapt its vocabulary to the testing and security domains. Besides, denoising the input test method to remove excessively long strings and version numbers could be beneficial. Concerning the Matching task, the errors were also due to the limited understanding of the vulnerability, caused by the too short description given by CVE. This could be mitigated by adding additional sources to fully comprehend the vulnerability, e.g., issue trackers and mailing lists.
@@
The Usefulness of Witnessing Tests. The release of Vul4J in 2022 paved the way for numerous software security tasks. The ability of VuTeCo to find security-related tests can help expand the known body of vulnerability-witnessing tests, enabling activities like the automated generation of security unit tests (Kang et al., 2022; Chen et al., 2024; Gao et al., 2025). Nevertheless, the potential applications of vulnerability-witnessing tests extend far beyond this (Pan et al., 2024). For example, the witnessing tests can act as proofs-of-concept supporting the automatic generation of realistic exploits, building on advances in security test generation models (Chen et al., 2024; Gao et al., 2025). Besides, they can support the automated vulnerability repair (AVR) process by localizing the vulnerable statements or assessing the plausibility of a generated patch (Mohammadi et al., 2019b; Ságodi et al., 2024; Zhou et al., 2024; Bui et al., 2024). We also envision using witnessing tests to support the retrieval of vulnerability-contributing commits (Bao et al., 2022; Hinrichs et al., 2025), as they can be run to triangulate when a vulnerability was introduced in a project; to the best of our knowledge, this task has not been investigated yet. Additionally, software engineers can benefit from access to a variety of example test cases. For instance, they can reuse tests from past vulnerabilities to address similar issues in their projects. The dataset Test4Vul, which we publicly release for the research community (Iannone, Emanuele and Bui, Quang-Cuong and Scandariato, Riccardo, 2025), is designed to achieve such foreseen applications.
@@
The Anatomy of Witnessing Tests. The retrieval of witnessing tests is challenging mainly due to the lack of empirical knowledge on what such tests look like. To date, no study has outlined a clear profile of tests that witness vulnerabilities or, more broadly, unit tests focused on security. The absence of characterization of vulnerability tests entails a significant knowledge gap, especially when compared with traditional functional tests. For instance, we are unaware of the setup required before calling the vulnerable component, what assertions should examine, or the number of tests required to “cover” all relevant scenarios for a vulnerability type. The only commonality between the two test types is that they both aim to identify undesirable behaviors in the code that violate certain requirements or properties. We believe this lack of knowledge might be ascribable to the difficulty in formulating security requirements at the unit/component level (i.e., methods or classes) since they are often considered at the system level (Mai et al., 2018; Felderer et al., 2016). Unfortunately, shedding light on these aspects requires numerous examples of witnessing tests. In fact, this study was conducted to address this shortage by providing an approach (VuTeCo) to expand the knowledge base of witnessing tests and to draw more attention to this topic. Once a line is drawn between vulnerability-witnessing tests and traditional tests, innovative solutions can be designed to help developers write more security tests.
@@ 8. Conclusion
We presented VuTeCo, a framework for finding security-related tests in Java projects and matching them to their corresponding vulnerabilities. The experimental evaluation (RQ1) identified promising AI models for Finding and Matching tasks, while the assessment in the wild (RQ2) demonstrated the usefulness of VuTeCo. The manually-confirmed tests have been collected in a novel dataset, Test4Vul, which we have publicly released. VuTeCo is the first solution explicitly addressing the problem of finding security unit tests in software repositories, laying the foundation for future research in this area.
@@
The paper’s replication package is available on FigShare (Iannone, Emanuele and Bui, Quang-Cuong and Scandariato, Riccardo, 2025). It contains the implementation of VuTeCo (as in this paper), the dataset Test4Vul, the scripts to reproduce the experiments, and the resulting raw and processed data.

### v4

@@ 4.1.1. Finding Task
Table 1 reports the performance of the six evaluated models and the three baseline approaches in their best configurations for the Finding task on $TE_{F}$ extracted from Vul4J (the other configurations are in the replication package (Iannone, Emanuele and Bui, Quang-Cuong and Scandariato, Riccardo, 2026)). The best model was UniXcoder (Guo et al., 2022), achieving an $F_{0.5}$ score of $0.73$, supported by a high precision of $0.83$. This means that it can often recognize real vulnerability-witnessing tests, with minimal false positives (just two). In other words, the tests that it flags are very likely to be security-related. The good performance of UniXcoder was also confirmed by the high MCC score of $0.61$, indicating a strong positive correlation with the ground truth (following the same interpretation of Pearson’s correlation coefficient (Pearson and Galton, 1895)). UniXcoder identified 10 out of 21 vulnerability-witnessing tests in $TE_{F}$, resulting in an unremarkable recall score of $0.48$. Still, this was the highest recall across all tested models. UniXcoder had slightly lower precision than Qwen2.5-Coder, which was $0.88$; however, this was due to a single false positive.
@@ 4.1.2. Matching Task
Table 2 reports the performance of the six evaluated models and the three baseline approaches in their best configurations for the Matching task on $TE_{M}$ extracted from Vul4J (the other configurations are in the replication package (Iannone, Emanuele and Bui, Quang-Cuong and Scandariato, Riccardo, 2026)). The best model was DeepSeek Coder (Guo et al., 2024), achieving an $F_{0.5}$ score of $0.65$ and a high precision of $0.75$. This means that it can often validate pairs of test cases and vulnerability descriptions, with minimal false positives (just three). In other words, the matches that it returns are very likely to be valid. This is also confirmed by the MCC of $0.57$, indicating a strong positive correlation with the ground truth (Pearson and Galton, 1895).
Unlike the Finding task, the best model belongs to the LLM group. This DeepSeek Coder model was trained in “full-train” mode (PT-FT), which pre-trained on $TR^{\prime}_{M}$ (the simplified Matching task) for eight epochs and then fine-tuned on $TR_{M}$ (the regular Matching task) for five epochs. Both training datasets were augmented with SPAT (Yu et al., 2022) (ran five times). This confirmed that the model benefited from the two-phase training. Indeed, comparing this version with the configuration that only performed fine-tuning (FT) on $TR_{M}$, the $F_{0.5}$ score went from $0.58$ to $0.65$, due to an improved precision (from $0.64$ to $0.75$) and the same recall. The other LLMs, by contrast, achieved their best performance with only fine-tuning (i.e., pre-training actually lowered their scores). The effectiveness of two-phase training appears to vary with the specific model, making it an impactful factor. Furthermore, all LLMs showed consistent benefits from data augmentation. This is likely due to the inability to employ a weighted loss function as we did with the CRMs.
@@
Among the baseline approaches, SimMatch performed poorly. The SimMatchCRM flavor achieved greater precision, but at the cost of reducing the detection rate almost to zero (only one true positive and two false positives). Despite its simplicity, GrepMatch (with two required hits) achieved higher precision, i.e., $0.50$. This method appears suitable as an initial, lightweight approach for quickly matching a few test cases with vulnerabilities (i.e., the “low-hanging fruit”); however, it also missed many cases ($0.14$ recall).
@@ 4.2.1. Finding Task
Based on the results of RQ1, VuTeCo uses UniXcoder for the Finding task. After processing 1,105,491 test cases (from a total of 427 projects), VuTeCo flagged 319 test methods as “Security”, found in 83 projects. Among these, the manual assessment confirmed that 224 covered security-related aspects—i.e., VuTeCo scored $0.70$ precision. The inspectors agreed on 294 ($92\%$) cases, with $0.81$ Cohen’s Kappa score, indicating a strong inter-rater agreement. They jointly reviewed the remaining 25 cases with conflicting judgments until consensus was reached.
The precision score in the wild is not much farther from the $0.83$ observed during the experiment in RQ1—indeed, a non-negligible drop was anticipated. Thus, the transfer of the Finder model in the wild mostly preserved its capability. We observed that VuTeCo was misled by test cases containing keywords commonly associated with vulnerabilities, even though these tests did not actually uncover any security issues. For instance, a test case in Apache Struts contains terms like ‘bad’, ‘pollution’, and ‘inject’—which are often found in security vocabularies—despite having a different meaning there (Apache Flink, 2025). Besides, we observed that test cases containing literal strings such as paths (e.g., URLs or file paths), shell commands, version numbers, or hashes are more likely to be incorrectly flagged. Such problems could be addressed by introducing fine-grained semantic analyses and improved keyword matching to reduce false positives.
@@ 4.2.2. Matching Task
Based on the results of RQ1, VuTeCo uses DeepSeek Coder for the Matching task. After processing 5,451,212 pairs of test cases and vulnerability descriptions, VuTeCo flagged 96 pairs as “Matched”, involving 35 unique tests from 19 projects. Among these, the manual assessment confirmed that 45 were correct, i.e., VuTeCo scored $0.47$ precision. The inspectors agreed on 85 ($89\%$) cases, with $0.77$ Cohen’s Kappa score, indicating a substantial inter-rater agreement. They jointly reviewed the remaining 11 cases with conflicting judgments until consensus was reached.
Unlike the Finding task, here we observed a larger drop in precision from the $0.75$ achieved in RQ1. Many of the invalid matches were due to the similarity between the terms in the test case and the CVE description (just as in the Finding case). This decline in precision can be attributed to two main reasons: (i) the 427 projects chosen for this assessment might have no valid match at all, whereas the 51 projects used in RQ1 were guaranteed to have at least one valid match; (ii) the number of vulnerabilities for each project is higher than in RQ1, particularly for large projects like Spring Framework or Jenkins. Therefore, there is a natural distribution shift that increases the likelihood of false positives. Despite these issues, the transfer of the Matcher model in the wild was deemed acceptable, though it requires further improvements, such as better prompting and safeguards against misleading terms.
In the end, the 35 tests correctly matched with the right vulnerability were added to the new dataset Test4Vul (Iannone, Emanuele and Bui, Quang-Cuong and Scandariato, Riccardo, 2026) alongside the 224 security-related tests from the Finding task, totaling 259.
Answer to RQ2. VuTeCo found 224 security-related test cases in the wild, i.e., $70\%$ of all tests returned. The false positives were primarily due to security-related terms and constant strings appearing in the test code. VuTeCo returned 45 correct matches, i.e., $47\%$ of the total matches. Test code and CVE descriptions with similar vocabulary were matched incorrectly.
@@ 5. Discussion
The Retrieval of Witnessing Tests. The main advantage of VuTeCo lies in its fully static nature. A dynamic assessment would require building an entire project and setting up the environment to execute all test cases, which may not always succeed due to dependency issues. The extent of the results that VuTeCo returned in the wild (RQ2) permits manual inspections to be done in a reasonable time. Indeed, VuTeCo greatly reduced the search space by four orders of magnitude, from millions to a few hundred cases. Hence, VuTeCo is designed as a lightweight assessment tool to find vulnerability-witnessing tests before proceeding to a more comprehensive dynamic assessment. Despite the positive results from the experimental evaluation (RQ1), retrieving vulnerability-witnessing tests remains a challenging task. The main limitation of the AI models we evaluated is their low recall, which did not exceed 0.48 on the Finding task and 0.43 on the Matching task. This may stem from the training data: Vul4J contains few, non-diverse security tests, which prevents the models from learning general patterns (for instance, Vul4J has only six XSS tests and two authentication tests). Beyond labeling more example tests (also a goal of this paper), the low recall could be mitigated by combining VuTeCo’s AI models with heuristic methods that achieve higher recall. Furthermore, we identified the main reasons for VuTeCo’s false-positive classifications, finding that they may be attributable to the test code’s vocabulary. Tests containing terms and constant strings often related to security led the AI models into error, likely because they were similar to the vocabulary of the positive instances in the training sets for the Finding and Matching tasks. This problem highlights the need for a dedicated pre-training session to adapt the model’s vocabulary to testing and security domains. Denoising the input test methods to remove excessively long strings and version numbers may also help. For the Matching task, errors may stem from a limited understanding of the vulnerability arising from overly short and vague CVE descriptions. This could be mitigated by incorporating additional vulnerability sources, such as issue trackers and mailing lists.
@@
The Usefulness of Witnessing Tests. The release of Vul4J in 2022 paved the way for numerous software security tasks. The ability of VuTeCo to find security-related tests helps expand the labeled body of known vulnerability-witnessing tests, enabling activities like the automated generation and improvement of security unit tests (Kang et al., 2022; Chen et al., 2024; Gao et al., 2025). Nevertheless, the potential applications of vulnerability-witnessing tests extend far beyond this (Pan et al., 2024). For example, the witnessing tests can act as proofs-of-concept supporting the automatic generation of realistic exploits (Chen et al., 2024; Gao et al., 2025). Besides, they can support the automated vulnerability repair (AVR) process by localizing the vulnerable statements or assessing the correctness of a generated patch (Mohammadi et al., 2019b; Ságodi et al., 2024; Zhou et al., 2024; Bui et al., 2024). Researchers can also use such examples to characterize security-related tests by distilling the recurring test structure for each vulnerability type. This would enable comprehending the typical setup actions and assertions involved (further details in the next discussion point). We also envision using witnessing tests to support the retrieval of vulnerability-contributing commits (Bao et al., 2022; Hinrichs et al., 2025), as they can be run to triangulate when a vulnerability was introduced in a project. Additionally, software developers can benefit from access to a variety of example test cases. For instance, they can reuse tests from past vulnerabilities to address similar issues in their projects. The dataset Test4Vul, which we publicly release for the research community (Iannone, Emanuele and Bui, Quang-Cuong and Scandariato, Riccardo, 2026), is designed to achieve such foreseen applications.
@@
The Anatomy of Witnessing Tests. The retrieval of witnessing tests is challenging, mainly because there is limited empirical knowledge about what such tests look like. To date, no study has outlined a clear profile of tests that witness vulnerabilities or, more broadly, unit tests focused on security. The lack of characterization of vulnerability tests entails a significant knowledge gap, especially when compared with traditional functional tests. For instance, we are unaware of the setup required before calling the vulnerable component, what assertions should examine, or the number of tests required to “cover” all relevant scenarios for a vulnerability type. The only commonality between the two test types is that they both aim to identify undesirable behaviors in the code that violate certain requirements or properties. We believe this lack of knowledge might be ascribable to the difficulty in formulating security requirements at the unit/component level (i.e., methods or classes) since they are often considered at the system level (Mai et al., 2018; Felderer et al., 2016). Unfortunately, shedding light on these aspects requires numerous examples of witnessing tests. In fact, this study was conducted to address this shortage by providing an approach (VuTeCo) to expand the knowledge base of witnessing tests and to draw more attention to this topic. Once a line is drawn between vulnerability-witnessing tests and traditional tests, innovative solutions can be designed to help developers write more security tests.
@@ 8. Conclusion
We presented VuTeCo, a framework for finding security-related tests in Java projects and matching them to their corresponding vulnerabilities. The experimental evaluation (RQ1) identified promising AI models for Finding and Matching tasks, while the assessment in the wild (RQ2) demonstrated the usefulness of VuTeCo. The manually-confirmed tests have been collected in a novel dataset, Test4Vul, which we have publicly released. VuTeCo is the first solution explicitly tackling the problem of finding security unit tests in software repositories, laying the foundation for future research.
@@
The active version of VuTeCo is available at: https://github.com/tuhh-softsec/vuteco. Test4Vul is available at: https://github.com/tuhh-softsec/test4vul. The paper’s replication package is available on Zenodo (Iannone, Emanuele and Bui, Quang-Cuong and Scandariato, Riccardo, 2026) and contains the implementation of VuTeCo used in this work, the dataset Test4Vul, the scripts to reproduce the experiments, and all raw and processed data.